OpenStudy (anonymous):

If f(x) = 4x - 1 and g(x) = x^2 + 2x + 2, what is f(g(x)) ? x2 + 8x - 2 x2 + 6x + 1 4x2 + 6x - 1 4x2 + 8x + 7

jimthompson5910 (jim_thompson5910):

f(x) = 4x-1 f(g(x)) = 4(g(x)) - 1 f(g(x)) = 4(x^2+2x+2) - 1 f(g(x)) = 4x^2+8x+8 - 1 f(g(x)) = 4x^2+8x+7